I know I have studied this all week but for the life of me I can not figure out how to do this. Please help.
Listed below are the durations(in hours)of a simple random sample of all flights of a space shuttle program. Find the range, variance, and standard deviation for the sample data. Is the lowest duration time unusual? Why or why not?
71 98 233 194 166 270 195 379 258 230 390 340 230 243 0
The range of the simple data is____ hours.
The variance of the sample data is_____.
The standard deviation of the sample data is ____ hours.
Is the lowest duration time unusual? Why or why not?

Answer by stanbon(75887)

Find the range, variance, and standard deviation for the sample data. Is the lowest duration time unusual? Why or why not?
71 98 233 194 166 270 195 379 258 230 390 340 230 243 0
The range of the simple data is highest-lowest = 390-0 = 390 hours.
Using my TI-84 I get the following::
The variance of the sample data is 107.94^2.
----------------------
The standard deviation of the sample data is 107.94 hours.
----------------------
Is the lowest duration time unusual?
Why or why not?
z(x) = (x-u)/s
z(0) = (0-219.8)/107.94 = -2.036
---
Unusual because it is more than 2 standard deviations below the mean.
===================
Cheers,
Stan H.